# Bayesian competing risks survival modeling for assessing the cause of death of patients with heart failure

### v1

@@ 4.1 Model specification
We return to the cardiac resynchronization therapy study, whose data we will analyse using a Bayesian competing model with two competing events, CV and N-CV death. Each patient undergoing CRT (initial state) can access any of them after having surgery. We define $T_{C}$ and $T_{NC}$ as the survival time associated with CV and N-CV death, respectively. We express the cause-specific hazard function for CV and N-CV death in terms of a Cox proportionals hazards modelling,
$h_{c}(t\mid\bm{\phi}_{c},\,\bm{\beta}_{c})=\,h_{0,c}(t\mid\bm{\phi}_{c})\,\exp\{\bm{x}^{\top}_{c}\bm{\beta}_{c}\},$
$h_{nc}(t\mid\bm{\phi}_{nc},\,\bm{\beta}_{nc})=\,h_{0,nc}(t\mid\bm{\phi}_{nc})\,\exp\{\bm{x}^{\top}_{nc}\,\bm{\beta}_{nc}\},$
@@
$\bm{x}^{\top}_{c}\bm{\beta}_{c}=\,\,\beta_{c.1}I_{Man}+\beta_{c.2}I_{Pace}+\beta_{c.3}I_{Ische}+\beta_{c.4}I_{III}+\beta_{c.5}I_{IV}$
@@
$\bm{x}^{\top}_{nc}\,\bm{\beta}_{nc}=\,\,\beta_{nc.1}I_{Man}+\beta_{nc.2}I_{Pace}+\beta_{nc.3}I_{Ische}+\beta_{nc.4}I_{III}+\beta_{nc.5}I_{IV}+$
@@
The term $I_{A}$ is an indicator function whose value is 1 if the subsequent individual has the characteristic $A$ and 0 otherwise. In particular, the five indicator functions in the two expressions above, $I_{Man}$, $I_{Pace}$, $I_{Ische}$, $I_{III}$, and $I_{IV}$ refer to being man, having a pacemaker implanted, ischaemic heart disease, and functional classification III and IV, respectively. Baseline covariates Age, Crea, LBBB, PAF, and QRS describe for each individual their age, level of creatinite, LBBB, previous AF, and QRS, respectively. The reference category for both causes of death is being woman, having an implanted ICD, dilated heart disease, and functional classification II.
@@
$h_{0.c}(t\mid\bm{\phi}_{c})=\lambda_{c}\,\alpha_{c}\,t^{\alpha_{c}-1},$
@@
$h_{0.c}(t\mid\bm{\phi}_{c})=\mbox{$\sum$}_{r=1}^{R_{c}}\,\gamma_{c.r}\,I_{{(a_{c.r-1},a_{c.r}]}},$
$h_{0.nc}(t\mid\bm{\phi}_{nc})=\mbox{$\sum$}_{r=1}^{R_{nc}}\,\gamma_{nc.r}\,I_{{(a_{nc.r-1},a_{nc.r}]}},$ (5)
@@
We need to complete the Bayesian model with the specification of a prior distribution for all the parameters of the model, $\bm{\theta}=(\bm{\phi}_{c},\bm{\phi}_{nc},\bm{\beta}_{c},\bm{\beta}_{nc})$. We assume prior independence and have selected, as much as possible, non-informative prior distributions that are standard in this application environment in order to give all the prominence of the inferential process to the data. We have elicited normal distributions $N(0,0.001)$ for the coefficients of regression in $\bm{\beta}_{c}$, $\bm{\beta}_{nc}$. In the case of Weibull baseline risk functions, we have elicited gamma distributions $\text{Ga}(0.01,0.01)$ for the $\lambda$’s and uniforms $\text{U}(0,10)$ for the $\alpha$’s. Alternatively, in the case of the risk functions defined through the piecewise functions in (5), we have selected gamma distributions $\text{Ga}(0.01,0.01)$ for the $\phi$’s coefficients. We are working with two models that are identical except for the specification of the baseline risk function. The posterior distribution $\pi(\bm{\theta}\mid\mathcal{D})$ of each of them has been approximated via Bayesian inference through MCMC with JAGS software (Plummer, 2003) through the rjags package (Plummer, 2023) in the R version 4.1 (R Core Team, 2023). Three MCMC chains, 100,000 iterations, 1,000 burning iterations and a thinning interval of 10 have been used to compute each model (Álvares et al., 2021). For each model, convergence has been checked via the Gelman-Rubin statistic.
Table LABEL:tab:2 includes a summary of the approximate posterior distribution of the parameters of both models (whose only formal difference is in the specification of the hazard function). In particular, the posterior mean and standard deviation of the coefficients of both models are presented. Regarding the results, we would first like to comment on the similarity between the estimates of the parameters associated with the different covariates in both models, both for CV and N.CV death. It is also interesting to note that in both types of death, the posterior standard deviation associated with the categorical covariates is greater than that estimated for the quantitative covariates and that greater uncertainty is observed in the Cr death group. In the case of CV death, the categories of being a man, have a pacemaker device, an ischaemic etiology of the disease, and a functional classification III and IV generate an increase in the hazard function and, therefore, lower survival times. Quantitative covariates also have a negative impact on survival, except for LBBB and certainly QRS. In the N-CV death group, it is worth noting the low relevance of gender and the type of device implanted. In addition, note that having a diagnosis in functional classification IV decreases the risk of death, but age increases this risk notably.
@@
We have not detected any relevant change in the posterior distribution obtained from these new prior distributions compared with the results presented in Table LABEL:tab:3. We therefore have results that we could describe as robust.
@@ 4.2 Model uncertainty
Table LABEL:tab:3 shows the five competing risk models that the García-Donato et al. (2023) procedure selects with the highest probability. We again confirm that the results obtained for both causes of death are very different. We have little uncertainty about the best model in the case of N-CV death, with only covariate Age; and a little more uncertainty in the case of CV death: Creatinine appears in all the models and, in fact, the model with the highest probability only includes this covariate. CRT Device and Etiology appear in three of the five models and Age in only one.
Table 3: Posterior model probabilities of the five most probable competing risk models
@@
The spike and slab method applies to all elements in the model, the baseline risk function and the covariates. We recall that we have worked with two different formulations of the basal risk function, one through the Weibull distribution and the second through a piecewise-defined function. A $N(0,0.001)$ prior distribution has been selected for modeling each slab term and the beta distribution Beta(1,1) has been elicited associated to probability $\eta$ in (4).
Table 4: Inclusion probabilities of covariates associated with cardiovascular and non-cardiovascular death determined through the Bayes factor and the spike and slab method, this latter with a Weibull and a piecewise baseline risk function.
@@
Table LABEL:tab:4 shows the estimated probability of inclusion in the model of each of the covariates provided by the Bayes factor method and by a spike distribution method, the latter with the Weibull and the piecewise baseline risk function. We can observe that the results obtained with the spike are much more drastic than with the procedure based on the Bayes factor because they only give a high inclusion probability estimate to the covariate Creatinine in the case of CV death and to Age in the case of N-CV death. Both covariates are also the favorites of the Bayes factor. No relevant differences were observed between the two modeling approaches based on the spike-and-slab method. The model with the highest posterior probability and the model that includes all covariates with inclusion probability 0.5 or higher (Median posterior probability model, Barbieri and Berger, 2004) coincide in this case: Creatinine for CV death and Age for N-CV death.
@@
As a final step, we used the DIC, WAIC and LPML criteria to compare four models: the full model with all covariates, the model that only includes covariates that have a probability of inclusion greater than 0.5, the model with covariates with inclusion probabilities greater than 0.4, and the previous model with the NYHA covariate added. In all models, we have considered a Weibull and a piecewise polynomial baseline risk function. All the models have a fairly similar behavior with the three criteria, although the models with Weibull baseline risk function are always better and also, in all cases, the last one is the best. Finally, we decided to select the latter model: Weibull baseline hazard function for all events and covariates. Age, Creatinine, Device, Etiology and NYHA for CV death, and Age for N-CV death.
@@ 4.3 Selected model and posterior outputs
Table LABEL:tab:5 shows a summary of the approximated posterior distribution for the parameters of the selected model. We have used JAGs in the same way as with the full model to obtain an approximate sample from the subsequent posterior distribution. If we observe the results of the complete model and the selected model, we can see stability and similarity in the information corresponding to the common covariates in both events, especially for CV death. In this case, the ischemic etiology is associated with an increased risk of death compared to the dilated etiology. This also occurs with the CRT pacemaker with regard to the defibrillator device. Age, Creatinine and NYHA functional classification III and mainly IV have also a positive impact on CV death. Age is the only relevant covariate for N-CV death, with a strong positive impact on the risk of death.
@@
In the case of CV death the survival of type A patients was less than 10 and 15 years with probabilities 0.24 and 0.29. These probabilities turn out to be 0.67 and 0.73 for type B patients. Adding the condition that the patient is still alive 3 years after implantation, these probabilities become 0.15 and 0.23 for type A patients and 0.54 and 0.62 for type B patients,. respectively. In the case of having lived more than 5 years after transplantation, the probabilities of living less than 10 years become 0.12 and 0.17 for type A patients, and 0.46 and 0.61 for type B patients. Finally, for patients who survive more than 8 years after implantation, the probability of a CV-death less than 10 or 15 years after implantation is 0.05 and 0.13 for type A patients and 0.24 and 0.50 for type B patients, respectively.
@@ 5 Conclusions
We have proposed a Bayesian approach to competing risk survival models through the so-called cause-specific hazard models which we have complemented through a multistate model perspective. We have carried out an inferential process with non-informative prior distributions and MCMC methods to approximate the respective posterior distribution. Outputs of interest in the statistical study also include the posterior distribution for the overall survival function of the time in which the first event occurs and the cumulative incidence function of that survival time due to each event-cause. In addition, we have discussed the posterior distribution for the transition probabilities between the different states of the multistate model both from the initial event and the case where the individual under consideration remains under the initial situation for a certain period of time.
Model uncertainty has been adressed for both variable selection and the specification of the baseline hazard function in the subsequents proportional hazard functions associated to the time in wchich the event occurs for each cause. The proposal by García-Donato et al (2023) with regard Bayes factor for univariate Cox regression is generalised to competing risk models. An extension of the CPO, which is the basis for logarithm of the pseudomarginal likelihood criterion based on the predictive power of the model, for competing risks scenarios is also presented.
@@
This research was supported by the Xunta de Galicia via an industrial doctoral grant to the first author, grants SBPLY/17/180501/000491 and SBPLY/21/180501/000241, funded by Consejería de Educación, Cultura y Deportes (JCCM, Spain) and Fondo Europeo de Desarrollo Regional, and the Ministerio de Ciencia, Innovación y Universidades through the project “PID2022-136455NB-I00 MCIN/AEI/10.13039/501100011033/FEDER, UE” and the European Regional Development Fund.
Conflict of Interest
The authors have declared no conflict of interest.
Data availability
The JAGS codes for reproducing the analyses are available on GitHub: https://github.com/jegubo/BUGS_Competing.git.

### v2

@@ 4.1 Model specification
We return to the cardiac resynchronization therapy study, whose data we will analyse using a Bayesian competing risks model with two competing events, CV and N-CV death. Each patient undergoing CRT (initial state) can access any of them after having surgery. We define $T_{C}$ and $T_{NC}$ as the survival time associated with CV and N-CV death, respectively. We express the cause-specific hazard function for CV and N-CV death in terms of a Cox proportionals hazards modelling,
$h_{c}(t\mid\boldsymbol{\phi}_{c},\,\boldsymbol{\beta}_{c})=\,h_{0,c}(t\mid\boldsymbol{\phi}_{c})\,\exp\{\boldsymbol{x}^{\top}_{c}\boldsymbol{\beta}_{c}\},$
$h_{nc}(t\mid\boldsymbol{\phi}_{nc},\,\boldsymbol{\beta}_{nc})=\,h_{0,nc}(t\mid\boldsymbol{\phi}_{nc})\,\exp\{\boldsymbol{x}^{\top}_{nc}\,\boldsymbol{\beta}_{nc}\},$
@@
$\boldsymbol{x}^{\top}_{c}\boldsymbol{\beta}_{c}=\,\,\beta_{c.1}I_{Man}+\beta_{c.2}I_{Pace}+\beta_{c.3}I_{Ische}+\beta_{c.4}I_{III}+\beta_{c.5}I_{IV}$
@@
$\boldsymbol{x}^{\top}_{nc}\,\boldsymbol{\beta}_{nc}=\,\,\beta_{nc.1}I_{Man}+\beta_{nc.2}I_{Pace}+\beta_{nc.3}I_{Ische}+\beta_{nc.4}I_{III}+\beta_{nc.5}I_{IV}+$
@@
Baseline covariates Age, Crea, LBBB, PAF, and QRS describe for each individual their age, level of creatinite, LBBB, previous AF, and QRS, respectively. The reference category for both causes of death is being woman, having an implanted ICD, dilated heart disease, and functional classification II.
@@
$h_{0.c}(t\mid\boldsymbol{\phi}_{c})=\lambda_{c}\,\alpha_{c}\,t^{\alpha_{c}-1},$
@@
$h_{0.c}(t\mid\boldsymbol{\phi}_{c})=\mbox{$\sum$}_{r=1}^{R_{c}}\,\gamma_{c.r}\,I_{{(a_{c.r-1},a_{c.r}]}},$
$h_{0.nc}(t\mid\boldsymbol{\phi}_{nc})=\mbox{$\sum$}_{r=1}^{R_{nc}}\,\gamma_{nc.r}\,I_{{(a_{nc.r-1},a_{nc.r}]}},$ (5)
@@
Sex (Man) 0.092 0.427 0.047 0.422
We need to complete the Bayesian model with the specification of a prior distribution for all the parameters of the model, $\boldsymbol{\theta}=(\boldsymbol{\phi}_{c},\boldsymbol{\phi}_{nc},\boldsymbol{\beta}_{c},\boldsymbol{\beta}_{nc})$. We assume prior independence and have selected, as much as possible, non-informative prior distributions that are standard in this application environment in order to give all the prominence of the inferential process to the data. We have elicited normal distributions $N(0,0.001)$ for the coefficients of regression in $\boldsymbol{\beta}_{c}$, $\boldsymbol{\beta}_{nc}$. In the case of Weibull baseline hazard functions, we have elicited gamma distributions $\text{Ga}(0.01,0.01)$ for the $\lambda$’s and uniforms $\text{U}(0,10)$ for the $\alpha$’s. Alternatively, in the case of the hazard functions defined through the piecewise functions in (5), we have selected gamma distributions $\text{Ga}(0.01,0.01)$ for the $\phi$’s coefficients. We are working with two models that are identical except for the specification of the baseline hazard function. The posterior distribution $\pi(\boldsymbol{\theta}\mid\mathcal{D})$ of each of them has been approximated via Bayesian inference through MCMC with JAGS software (Plummer, 2003) through the rjags package (Plummer, 2023) in the R version 4.1 (R Core Team, 2023). Three MCMC chains, 100,000 iterations, 1,000 burning iterations and a thinning interval of 10 have been used to compute each model (Álvares et al., 2021). For each model, convergence has been checked via the Gelman-Rubin statistic.
Table LABEL:tab:2 includes a summary of the approximate posterior distribution of the parameters of both models (whose only formal difference is in the specification of the hazard function). In particular, the posterior mean and standard deviation of the coefficients of both models are presented. Regarding the results, we would first like to comment on the similarity between the estimates of the parameters associated with the different covariates in both models, both for CV and N-CV death. It is also interesting to note that in both types of death, the posterior standard deviation associated with the categorical covariates is greater than that estimated for the quantitative covariates and that greater uncertainty is observed in the Cr death group. In the case of CV death, the categories of being a man, have a pacemaker device, an ischaemic etiology of the disease, and a functional classification III and IV generate an increase in the hazard function and, therefore, lower survival times. Quantitative covariates also have a negative impact on survival, except for LBBB and certainly QRS. In the N-CV death group, it is worth noting the low relevance of gender and the type of device implanted. In addition, note that having a diagnosis in functional classification IV decreases the risk of death, but age increases this risk notably.
@@
We have not detected any relevant change in the posterior distribution obtained from these new prior distributions compared with the results presented in Table LABEL:tab:2. Furthermore, fitting metrics (DIC, WAIC and LPML) have been obtained for all the models considered, yielding similar results. In the light of the robustness of the posterior distribution obtained, we decided to keep the simpler and more ”standard” prior distribution for the regression coefficients: a $N(0,0.001)$ prior. Regarding the specification for the baseline piecewise function, and as it can be checked in the following sections, choosing a Weibull distribution instead always improves the model fit.
@@ 4.2 Model uncertainty
Table LABEL:tab:3 shows the five competing risks models that the García-Donato et al. (2023) procedure selects with the highest probability. We again confirm that the results obtained for both causes of death are very different. We have little uncertainty about the best model in the case of N-CV death, with only covariate Age; and a little more uncertainty in the case of CV death: Creatinine appears in all the models and, in fact, the model with the highest probability only includes this covariate. CRT Device and Etiology appear in three of the five models and Age in only one.
Table 3: Posterior model probabilities of the five most probable competing risks models
@@
The spike and slab method applies to all elements in the model, the baseline hazard function and the covariates. We recall that we have worked with two different formulations of the baseline hazard function, one through the Weibull distribution and the second through a piecewise-defined function. A $N(0,0.001)$ prior distribution has been selected for modeling each slab term and the beta distribution Beta(1,1) has been elicited associated to probability $\eta$ in (4).
Table 4: Inclusion probabilities of covariates associated with cardiovascular and non-cardiovascular death determined through the Bayes factor and the spike and slab method, this latter with a Weibull and a piecewise baseline hazard function.
@@
Sex 0.216 0.077 0.048 0.012 0.055 0.022
Table LABEL:tab:4 shows the estimated probability of inclusion in the model of each of the covariates provided by the Bayes factor method and by a spike distribution method, the latter with the Weibull and the piecewise baseline hazard function. We can observe that the results obtained with the spike are much more drastic than with the procedure based on the Bayes factor because they only give a high inclusion probability estimate to the covariate Creatinine in the case of CV death and to Age in the case of N-CV death. Both covariates are also the favorites of the Bayes factor. No relevant differences were observed between the two modeling approaches based on the spike-and-slab method. The model with the highest posterior probability and the model that includes all covariates with inclusion probability 0.5 or higher (Median posterior probability model, Barbieri and Berger, 2004) coincide in this case: Creatinine for CV death and Age for N-CV death.
@@
As a final step, we used the DIC, WAIC and LPML criteria to compare four models: the full model with all covariates, the model that only includes covariates that have a probability of inclusion greater than 0.5, the model with covariates with inclusion probabilities greater than 0.4, and the previous model with the NYHA covariate added. In all models, we have considered a Weibull and a piecewise constant baseline hazard function. All the models have a fairly similar behavior with the three criteria, although the models with Weibull baseline hazard function are always better and also, in all cases, the last one is the best. Finally, we decided to select the latter model: Weibull baseline hazard function for all events and covariates. Age, Creatinine, Device, Etiology and NYHA for CV death, and Age for N-CV death. This one is the model recommended by the cardiologist and the one with slightly better fitting metrics, as it can be seen in Table LABEL:tab:5.
Table 5: Fitting metrics (DIC, WAIC and LPML) for all the models considered. M1 refers to creatinine for CV death and age for NCV death. M2 refers to creatinine, etiology, DAI and age for CV death and age for NCV. M3 refers to creatinine, etiology, DAI, NYHA and age for CV death and age for NCV.
DIC WAIC LPML
Weibull baseline.
Complete 59200698 59200726 -367.97
M1 59200711 59200718 -359.25
M2 59200700 59200710 -355.44
M3 59200696 59200709 -355.25
Piecewise constant baseline.
Complete 59200917 59200934 -469.61
M1 59200931 59200937 -470.94
M2 59200920 59200929 -465.56
M3 59200915 59200925 -464.06
@@ 4.3 Selected model and posterior outputs
Table LABEL:tab:6 shows a summary of the approximated posterior distribution for the parameters of the selected model. We have used JAGS in the same way as with the full model to obtain an approximate sample from the subsequent posterior distribution. If we observe the results of the complete model and the selected model, we can see stability and similarity in the information corresponding to the common covariates in both events, especially for CV death. In this case, the ischemic etiology is associated with an increased risk of death compared to the dilated etiology. This also occurs with the CRT pacemaker with regard to the defibrillator device. Age, Creatinine and NYHA functional classification III and mainly IV have also a positive impact on CV death. Age is the only relevant covariate for N-CV death, with a strong positive impact on the risk of death.
@@
In the case of CV death the survival of type A patients was less than 10 and 15 years with probabilities 0.24 and 0.29. These probabilities turn out to be 0.67 and 0.73 for type B patients. Adding the condition that the patient is still alive 3 years after implantation, these probabilities become 0.15 and 0.23 for type A patients and 0.54 and 0.62 for type B patients. respectively. In the case of having lived more than 5 years after transplantation, the probabilities of living less than 10 years become 0.12 and 0.17 for type A patients, and 0.46 and 0.61 for type B patients. Finally, for patients who survive more than 8 years after implantation, the probability of a CV-death less than 10 or 15 years after implantation is 0.05 and 0.13 for type A patients and 0.24 and 0.50 for type B patients, respectively.
@@ 5 Conclusions
We have proposed a Bayesian approach to competing risks survival models through the so-called cause-specific hazard models which we have complemented through a multistate model perspective. We have carried out an inferential process with non-informative prior distributions and MCMC methods to approximate the respective posterior distribution. Outputs of interest in the statistical study also include the posterior distribution for the overall survival function of the time in which the first event occurs and the cumulative incidence function of that survival time due to each event-cause. In addition, we have discussed the posterior distribution for the transition probabilities between the different states of the multistate model both from the initial event and the case where the individual under consideration remains under the initial situation for a certain period of time.
Model uncertainty has been adressed for both variable selection and the specification of the baseline hazard function in the subsequents proportional hazard functions associated to the time in which the event occurs for each cause. The proposal by García-Donato et al (2023) with regard Bayes factor for univariate Cox regression is generalised to competing risks models. An extension of the CPO, which is the basis for logarithm of the pseudomarginal likelihood criterion based on the predictive power of the model, for competing risks scenarios is also presented.
@@ Acknowledgement
This research was supported by the Xunta de Galicia via an industrial doctoral grant to the first author (Doutoramento Industrial 2021, IN606D), and the Ministerio de Ciencia, Innovación y Universidades through the project
“PID2022-136455NB-I00 MCIN/AEI/10.13039/501100011033/FEDER, UE” and the European Regional Development Fund.